OpenStudy (anonymous):

Jonah was keeping track of how many cars drove by every hour. He realized that the number of cars were in a sequence that followed the pattern n + 4. How many cars drove by on the 12th hour? 48 8 16 3

OpenStudy (anonymous):

I assume n is the hour. 1+4=5 2+4=6 3+4=7 ..... 12+4=16.